. Below please find the production possibilities frontiers (PPFs) of Ghana and Spain.

Ghana PPF                                                                                        Spain PPF

Possible Combinations

Possible Combinations

Cars

T-shirts

Cars

T-shirts

0

8,000

0

12,000

100

6,000

100

9,000

200

4,000

200

6,000

300

2,000

300

3,000

400

0

400

0

  1. Given these PPFs, which country would you predict would specialize in each good? Why?
  2. Suppose a free trade deal is negotiated. What is the range of potential terms of trade?
Answer #1

A) Using all of it's resources, Ghana can produce either 8000 t-shirts or 400 cars. Therefore, Ghana's opportunity cost of producing one t-shirt is (400/8000) = (1/20) car and opportunity cost of producing 1 car is (8000/400) = 20 t-shirts.

On the other hand, using all of it's resources, Spain can produce either 12000 t-shirts or 400 cars. Therefore, Spain's opportunity cost of producing 1 t-shirt is (400/12000) = (1/30) car and opportunity cost of producing 1 car is (12000/400) = 30 t-shirts.

As, Ghana's opportunity cost of producing 1 car is less than that of Spain's, therefore Ghana has comparative advantage in production of car. And, Spain's opportunity cost of producing 1 t-shirt is less than that of Ghana's, so Spain has comparative advantage in production of t-shirt.

Therefore, Ghana will specialize in production of car and Spain will specialize in t-shirt production.

B) Both country will gain from trade if the price of trade lies between both countries' opportunity cost of production. Therefore, potential terms of trade would be (price of car in terms of t-shirts) 20 t-shirts < 1 car < 30 shirts. Or, alternatively, price of 1 t-shirt in terms of car should lie between (1/30) car and (1/20) car.